OpenStudy (mariebrittany07):

how do you solve 4x=3y-7 when x=4,5,6 i got x= 2.8 x=0.5 x=1.8 am i right can some one please help me

OpenStudy (anonymous):

Looks like that is wrong

OpenStudy (anonymous):

@mariebrittany07 u there?

OpenStudy (anonymous):

U need to substitute x=4 in 4x=3y-7 and solve for y.

OpenStudy (anonymous):

Can u substitute x=4 in 4x=3y-7

OpenStudy (anonymous):

It is like this: 4*4=3y-7

OpenStudy (anonymous):

Just plug in the value of x=4 in x

OpenStudy (anonymous):

First solve this: 4*4=3y-7

OpenStudy (anonymous):

Yes...... Now PLZ continue

OpenStudy (anonymous):

ADD 7 on both sides

OpenStudy (anonymous):

4*4=3y-7 16=3y-7 ADD 7 on both sides 16+7=3y-7+7

OpenStudy (anonymous):

A told u to ADD 7 on both sides

OpenStudy (anonymous):

-7+7=0

OpenStudy (anonymous):